Pearce’s Cricket Farm issued a 25-year, 6% semiannual bond 2 years ago. The bond currently sells for 92% of its face value. The company’s tax rate is 40%.

Suppose the book value of the debt issue is $45 million. In addition, the company has a second debt issue on the market, a zero coupon bond with 12 years left to maturity; the book value of this issue is $45 million and the bonds sell for 53% of par. Assume the par value of the bond is $1,000.

What is the company’s total book value of debt? (Enter the answer in dollars. Omit $ sign in your response.)

Total book value ?

What is the company’s total market value of debt? (Enter the answer in dollars. Omit $ sign in your response.)

Total market value           ?

What is your best estimate of the after-tax cost of debt? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places.)

Cost of debt ?
Answer #1

Book Value of Debt = 45 million + 45 Million = 90

b. Market Value of Debt = 92%*45+53%*45 = 65.25

c. YTM of first bond is calculated as follows

Let's assume par value of single bond = 1000
Coupon= 6%*1000/2 = 30
Price = 92%*1000 = 920
Number of Periods = (25-2)*2 = 46
YTM using excel formula =2*RATE(46,30,-920,1000) = 6.6860%


Price of second bond = 53%*1000 = 530
Number of periods = 12
YTM of second bond = (1000/530)^(1/12)-1 =5.4431%

Cost of debt = Cost of Debt * Weight of market value of debt + Cost of zero coupon debt* Weight of market value of zero coupon =6.6860%*92%*45/65.25+5.4431%*53%*45/65.25 =6.2317% or 6.23%

After tax cost of Debt = 6.2314%*(1-Tax rate) = 6.2317%*(1-40%) = 3.74%